Hilbert schemes and integrable systems

Kevin Costello

The cohomology ring of the Hilbert scheme of points on a smooth projective surface is described as a completely integrable quantum-mechanical system, which is an infinite particle limit of the Calogero-Sutherland model of n particles on a circle.